OpenStudy (anonymous):

an object is hanging by a string from your rear view mirror. while you are decelerating at a constant rate from 25 m/s to rest in 6.0 s, a) what angle does the string make with the vertical, and b) is it toward the windshield or away from it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do u no how to solve it

OpenStudy (agent0smith):

Find the rate of acceleration using: decelerating at a constant rate from 25 m/s to rest in 6.0 s Then draw a free body diagram, remember the net horizontal force on the object is due to the above acceleration.

OpenStudy (anonymous):

can u show me how

OpenStudy (lastdaywork):

@besmir13 Do you know the concept of relative motion ??

OpenStudy (anonymous):

actually no cus the professor just explained the chapter

OpenStudy (anonymous):

acceleration will be (final velocity-initial velocity)/6 seconds, horizontally gravitational acceleration is -9.8 m/s^2, vertically net acceleration vector will be sum of these two vectors

OpenStudy (agent0smith):

Tension is T, acceleration is a from above, angle is x vertical forces must be balanced so net force is zero, so: mg = Tcos(x) (the Tcos(x) is the vertical component of tension) Horizontal force is provided by the horizontal component of tension: net F = Tsin(x) ma = Tsin(x) Now solve the two equations (use substitution, solve one of them for T): mg = Tcos(x) ..... solve for T = mg/cos(x) ma = Tsin(x) You should get: tan(x) = a/g Solve for angle x.